OpenStudy (anonymous):

find values of b such that the vectors (-11,b,2) and (b,b^2,b) are perpendicular

OpenStudy (anonymous):

-11b+b^3+2b = 0 b^3 = 9b b^2= 9 b=+/- 3

OpenStudy (anonymous):

how do u know when its going to be perpendicular?

myininaya (myininaya):

dot product=0

myininaya (myininaya):

\[b^3-9b=b(b^2-9)=0 => b=0, b=3, or b=-3\]

OpenStudy (anonymous):

i for got about the 0

myininaya (myininaya):

:)

OpenStudy (anonymous):

thanks ya'll :)